# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ 3.2 Effective coherent state
$\|\eta_{s}\|^{2}$ $=\frac{\pi\lambda^{2}}{\kappa^{2}}\sum_{k\in\Gamma^{\textnormal{nor}}}\frac{\hat{V}(k)^{2}}{|k|}\left\{\log(2\kappa k_{\textnormal{F}}|k|s)-\textnormal{Ci}(2\kappa k_{\textnormal{F}}|k|s)+\gamma\right\}$
×$\times\left\{1+g(\kappa k_{\textnormal{F}}|k|s)\ \mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$
where $g:\mathbb{R}\to\mathbb{R}_{\geq 0}$ is a function independent of $k_{\textnormal{F}}$ and monotonically increasing.
Furthermore, define $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ by $(y,x)\mapsto f_{y}(x)\coloneqq\min\{e^{byx},(4x+2)^{ay}\}$ with $a\coloneqq\sqrt{2\pi}\|\hat{V}\|_{2},b\coloneqq\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}$. Then there exists a $C>0$ independent of $k_{\textnormal{F}}$ such that for $\lambda=\kappa,c_{0}>0$ and $k_{\textnormal{F}}$ sufficiently large
$\|\eta_{s}\|$ $\leq\log\left(f_{1}(\kappa k_{\textnormal{F}}s)\right),$
$e^{c_{0}\|\eta_{s}\|}$ $\leq f_{c_{0}}(\kappa k_{\textnormal{F}}s).$
@@
Note that $f_{y}$ is for all $y\geq 0$ monotonically increasing with $f_{y}(0)=1$. The same bounds hold also for $\eta_{s/\hbar}$ in the semiclassical regime with $\lambda=k_{\textnormal{F}}^{-2}$.
Proof.
Recall that by definition it holds
$\|\eta_{s}\|^{2}=\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|\big{(}\eta_{s}\big{)}_{\alpha}(k)|^{2}=\lambda^{2}\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|^{2}\underbrace{\sum_{\alpha\in\mathcal{I}_{k}}\bigg{|}n_{\alpha}(k)\frac{\sin\big{(}\epsilon_{\alpha}(k)s/2\big{)}}{\epsilon_{\alpha}(k)/2}\bigg{|}^{2}}_{\eqqcolon S_{k}}$ (5.1)
with $\epsilon_{\alpha}(k)=2\kappa k_{F}|k\cdot\omega_{\alpha}|$. We will first approximate $S_{k}$ and then give a upper bound.
$S_{k}$ $=\frac{1}{\kappa^{2}|k|}\int_{0}^{2\pi}\textnormal{d}\varphi\int_{0}^{\pi/2}\textnormal{d}\theta\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\sin\theta+\mathcal{E}$
$=\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(\kappa k_{\textnormal{F}}|k|su)^{2}}{u}+\mathcal{E}$
$=\frac{\pi}{\kappa^{2}|k|}\left\{\log(2\kappa k_{\textnormal{F}}|k|s)-\text{Ci}(2\kappa k_{\textnormal{F}}|k|s)+\gamma\right\}+\mathcal{E}$ (5.3)
$\mathcal{E}_{2}$ $=\left|\int_{p_{\alpha}}\textnormal{d}\sigma\ \frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}-\sigma(p_{\alpha})\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta_{\alpha})^{2}}{\cos\theta_{\alpha}}\right|$
$\leq\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\textnormal{d}}{\textnormal{d}\theta}\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\sup_{(\theta,\varphi)\in p_{\alpha}}\left|\theta-\theta_{\alpha}\right|\sigma(p_{\alpha})$
$\leq C(\kappa k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}$ (5.4)
$\mathcal{E}_{3}=\sup_{\hat{\omega}(\theta,\varphi)\in p_{\alpha}}\left|\frac{\sin(\kappa k_{\textnormal{F}}|k|s\cos\theta)^{2}}{\cos\theta}\right|\left|\sigma\Big{(}\bigcup_{\alpha\in\mathcal{I}_{k}}p_{\alpha}\Big{)}-\sigma(S)\right|\leq C\kappa k_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right).$ (5.5)
Thus with the triangle inequality
$\left|S_{k}-\frac{\pi}{\kappa^{2}|k|}\left\{\log(2\kappa k_{\textnormal{F}}|k|s)-\text{Ci}(2\kappa k_{\textnormal{F}}|k|s)+\gamma\right\}\right|$
$\leq\frac{C}{\kappa^{2}|k|}\left\{M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+(\kappa k_{\textnormal{F}}|k|s)^{2}M^{-\frac{3}{2}}+\kappa k_{\textnormal{F}}|k|s\left(N^{-\delta}+M^{\frac{1}{2}}N^{-\frac{1}{3}}\right)\right\}.$ (5.6)
We will now give an estimate for $S_{k}$ by approximating
$\frac{\sin(\alpha x)^{2}}{x}=\frac{1}{2}\frac{1-\cos(2\alpha x)}{x}\leq\begin{cases}\alpha^{2}x&,2\alpha x\in[0,\pi/2]\\ \frac{1}{x}&,2\alpha x>\pi/2\end{cases}$
where we used $1-x^{2}/2\leq\cos(x)$ for all $x\in[0,\pi/2]$. Therefore
$\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\frac{\sin(\kappa k_{\textnormal{F}}|k|su)^{2}}{u}$
$\leq\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(\kappa k_{\textnormal{F}}|k|s)^{2}u\ \chi(2\kappa k_{\textnormal{F}}|k|su\leq\pi/2)+\frac{1}{u}\chi(2\kappa k_{\textnormal{F}}|k|su>\pi/2)\right\}$
$\leq\frac{2\pi}{\kappa^{2}|k|}\int_{0}^{1}\textnormal{d}u\ \left\{(\kappa k_{\textnormal{F}}|k|s)^{2}u\ \chi\Big{(}u\leq\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}+\frac{1}{u}\ \chi\Big{(}u>\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}\right\}$
$=\frac{2\pi}{\kappa^{2}|k|}\left\{\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2}\Big{(}\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}^{2}-\log\Big{(}\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}\right\}\ \chi\Big{(}1>\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}$
$\qquad+\frac{2\pi}{\kappa^{2}|k|}\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}1\leq\frac{\pi}{4\kappa k_{\textnormal{F}}|k|s}\Big{)}$
$=\frac{2\pi}{\kappa^{2}|k|}\left\{\frac{\pi^{2}}{32}-\log(\pi)+\log(4\kappa k_{\textnormal{F}}|k|s)\right\}\ \chi\Big{(}\kappa k_{\textnormal{F}}|k|s>\frac{\pi}{4}\Big{)}$
$\qquad+\frac{2\pi}{\kappa^{2}|k|}\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2}\ \chi\Big{(}\kappa k_{\textnormal{F}}|k|s\leq\frac{\pi}{4}\Big{)}$
$\leq\frac{2\pi}{\kappa^{2}|k|}\min\left\{\frac{1}{2}(\kappa k_{\textnormal{F}}|k|s)^{2},\log(4\kappa k_{\textnormal{F}}|k|s+1)\right\}.$ (5.7)
Note that $\log(4\kappa k_{\textnormal{F}}|k|s+1)\leq\log(4\kappa k_{\textnormal{F}}s+1)+\log|k|$ using $\frac{1}{2}x^{2}\leq\log(4x+1)$ for all $x\leq 2$ and $|k|\geq 1$ for all $k\in\mathbb{Z}_{\ast}^{3}$. Thus, for $\lambda=\kappa$
$\|\eta_{s}\|^{2}$ $\leq 2\pi\sum_{k\in\mathbb{Z}_{\ast}^{3}}|\hat{V}(k)|^{2}\min\left\{\frac{(\kappa k_{\textnormal{F}}s)^{2}}{2}|k|,\frac{\log(4\kappa k_{\textnormal{F}}|k|s+1)}{|k|}\right\}$
$\leq 2\pi\min\left\{\|\hat{V}(\cdot)^{1/2}\|_{2}^{2}(\lambda k_{\textnormal{F}}s)^{2}/2,\|\hat{V}\|_{2}^{2}\left(\log(4\kappa k_{\textnormal{F}}s+1)+1\right)\right\}.$ (5.8)
$\|\eta_{s}\|$ $\leq\min\left\{b\kappa k_{\textnormal{F}}s,\log(4\kappa k_{\textnormal{F}}s+2)^{a}+a\right\},$ (5.9)
$e^{c_{0}\|\eta_{s}\|}$ $\leq\min\left\{e^{c_{0}b\kappa k_{\textnormal{F}}s},e^{c_{0}a}(4\kappa k_{\textnormal{F}}s+2)^{c_{0}a}\right\}$ (5.10)
with $a\coloneqq\sqrt{2\pi}\|\hat{V}\|_{2},b\coloneqq\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}.$
@@
It holds for all $s\in\mathbb{R}$, $n\in\mathbb{N}$, $\psi\in\mathcal{F}$
@@
Proof.
For the inequalities, we observe that for $n\in\mathbb{N}_{0}$ it holds
$\sum_{k\in\Gamma^{\textnormal{nor}}}\||k|^{n}\eta_{s}(k)\|_{l^{2}}$ $\equiv\sum_{k\in\Gamma^{\textnormal{nor}}}|k|^{n}\Big{(}\sum_{\alpha\in\mathcal{I}_{k}}|(\eta_{s})_{\alpha}(k)|^{2}\Big{)}^{1/2}=\lambda\sum_{k\in\Gamma^{\textnormal{nor}}}|k|^{n}|\hat{V}(k)|\left(S_{k}\right)^{1/2}.$ (5.11)
Since by assumption $\|(\cdot)^{n}\hat{V}\|_{1}$ is bounded for each $n\in\mathbb{N}$, the first statement follows. The second statement simply follows from the same calculation and recalling that $\langle\eta_{s},|k|^{n}\eta_{s}\rangle\equiv\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}|\eta_{\alpha}(k)|^{2}$. The third statement follows from Lemma A.4 and
$\|c^{\ast}(|k|^{n}\eta_{s})\psi\|\leq\sum_{k\in\Gamma^{\textnormal{nor}}}\|\sum_{\alpha\in\mathcal{I}_{k}}|k|^{n}\eta_{\alpha}(k)c_{\alpha}^{\ast}(k)\psi\|\leq\sum_{k\in\Gamma^{\textnormal{nor}}}\||k|^{n}\eta_{s}(k)\|_{l^{2}}\|(\mathcal{N}+1)^{1/2}\psi\|.$ (5.12)
@@
We use the approach as sketched in 3.8. Since the bosonic property only holds only with an error, the equality (3.17) holds only approximately:
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|\leq\int_{0}^{t}\textnormal{d}s\|h_{0}W(\eta_{s})\phi\otimes\Omega\|+\text{Error}_{1}+\text{Error}_{2}.$ (5.13)
@@
$\text{Error}_{1}\coloneqq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\left\{\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right\}W(\eta_{s})\mathcal{R}_{\alpha}(k)\psi.$ (5.14)
@@
$\text{Error}_{2}\coloneqq-$ $2i\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})\psi$
$\equiv$ $2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau e^{(1-\tau)B}\bigg{\{}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}e^{is\epsilon_{\alpha}(k)}\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)$
$+\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{-is\epsilon_{\alpha}(k)}(h_{y})_{\alpha}(k)\bigg{\}}e^{\tau B}\psi.$ (5.15)
Firstly, we show that the term $\mathcal{R}^{\lambda}_{\alpha}(k)\psi=\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi$ as defined in Lemma 2.3 constitutes indeed a small error. We estimate
$\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{\gamma\in\mathcal{I}_{l}}\|\mathcal{R}_{\gamma}(l)\psi\|^{2}$
$\leq\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{\gamma\in\mathcal{I}_{l}\cap\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}|\eta_{\gamma}(k)|\int_{0}^{1}\textnormal{d}\tau\|\mathcal{E}_{\gamma}(k,l)e^{\tau B}\psi\|\bigg{)}^{2}$
@@
where we used $\sum_{\alpha\in\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(k)|\bigg{)}^{2}\leq\sum_{k^{\prime},k\in\Gamma^{\textnormal{nor}}}\|\eta(k)\|_{l^{2}}\|\eta(k^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.1 and
@@
$\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\leq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\|c_{\alpha}^{\ast}(k)e^{(1-\tau)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|c_{\alpha}^{\ast}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
$\quad+\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|[c_{\alpha}^{\ast}(k),e^{(1-\tau)B}]\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
@@
$\quad+\sum_{k\in\Gamma^{\textnormal{nor}}}\|\eta_{s}\|\int_{0}^{1}\textnormal{d}\tau\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|\mathcal{R}_{\alpha}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}$
@@
where we used $[c_{\alpha}^{\ast}(k),e^{\lambda B}]=-\lambda\eta_{\alpha}(k)e^{\lambda B}-e^{\lambda B}\mathcal{R}_{\alpha}(k)$ from Lemma 2.3, the Cauchy-Schwarz inequality for the $\alpha$-summation, Lemma A.3 in the third inequality and in the fourth inequality we used Proposition 2.6 and (5.16).
@@
$\|\text{Error}_{1}\|$ $\leq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\left\{\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right\}e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\leq C\kappa k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\quad+C\lambda\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\|n_{\alpha}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\leq C\kappa k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}_{\alpha}(k)\psi\|$
$\quad+C\lambda k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\|\hat{V}\|_{2}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}_{\alpha}(k)\psi\|^{2}\bigg{)}^{1/2}$
$\leq C\kappa k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\ (e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{2}\psi\|$
@@
$\leq C\ (f_{C}(\kappa k_{\textnormal{F}}t)-1)(\lambda+\kappa)k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)^{2}\psi\|.$ (5.21)
where we used (A.2) and $e^{B}$ unitary in the third inequality and Lemma 5.1 in the last line. Using the choices of $\kappa=\lambda$ and $\psi=\phi\otimes\Omega$ we obtain the desired bound.

### v2

@@ 3.2 Effective coherent state
$\|\eta_{s}\|^{2}$ $=\pi\lambda^{2}\sum_{k\in\Gamma}\frac{\hat{V}(k)^{2}}{|k|}\left(\log(2k_{\textnormal{F}}|k|s)-\textnormal{Ci}(2k_{\textnormal{F}}|k|s)+\gamma\right)$
×$\times\left\{1+g(k_{\textnormal{F}}|k|s)\ \mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$
where $\gamma$ is the Euler-Mascheroni constant and $g:\mathbb{R}\to\mathbb{R}_{\geq 0}$ is a function independent of $k_{\textnormal{F}}$ and monotonically increasing.
Furthermore, define $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ by
$(y,x)\mapsto f_{y}(x)\coloneqq\min\{e^{\sqrt{\pi}\|\hat{V}(\cdot)^{1/2}\|_{2}yx},e^{\sqrt{2\pi}\|\hat{V}\|_{2}(\log(18)+\frac{1}{9})y}e^{\frac{\sqrt{8\pi}}{9}\|\hat{V}\|_{2}yx}\}.$
Then there exists a $C>0$ independent of $k_{\textnormal{F}}$ such that for $c_{0}>0$ and $k_{\textnormal{F}}$ sufficiently large
$\|\eta_{s}\|$ $\leq\log\left(f_{1}(\lambda k_{\textnormal{F}}s)\right),$ (5.25)
$e^{c_{0}\|\eta_{s}\|}$ $\leq f_{c_{0}}(\lambda k_{\textnormal{F}}s).$ (5.26)
@@
Note that $f_{y}$ is for all $y\geq 0$ monotonically increasing with $f_{y}(0)=1$.
The previous statement is useful when combined with the following estimate:
@@
There exists a constant $C>0$ only depending on $V$ such that it holds for all $s\in\mathbb{R}$, $n\in\mathbb{N}$, $\psi\in\mathcal{F}$
@@
We use the approach as sketched in Remark 3.9. Since the bosonic property holds only with an error, the equality (3.21) holds only approximately:
$\|e^{-i\mathbb{H}^{\text{eff}}t}\psi-e^{iP(t)}W(\eta_{t})\psi\|$ $=\|\psi-e^{i\mathbb{H}^{\text{eff}}t}e^{-iE_{N}^{\text{pw}}t}e^{i2\text{Im}(\nu_{t})}e^{-i\text{Im}\int_{0}^{t}\textnormal{d}s\langle\dot{\eta}_{s},\eta_{s}\rangle}W(\eta_{t})\psi\|$
$\leq\int_{0}^{t}\textnormal{d}s\|\big{(}\mathbb{H}^{\text{eff}}-E_{N}^{\text{pw}}+2\text{Im}(\dot{\nu}_{s})-\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle\big{)}W(\eta_{s})\psi-i\partial_{s}W(\eta_{s})\psi\|$
$\leq\int_{0}^{t}\textnormal{d}s\|h_{0}W(\eta_{s})\phi\otimes\Omega\|+\|\text{Error}_{1}\|+\|\text{Error}_{2}\|.$ (5.27)
@@
$\text{Error}_{1}\coloneqq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\left(\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right)W(\eta_{s})\mathcal{R}^{1}_{\alpha}(k)\psi.$ (5.28)
@@
$\text{Error}_{2}\coloneqq-$ $2i\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\psi$
$\equiv$ $2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau e^{(1-\tau)B}\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}e^{is\epsilon_{\alpha}(k)}\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)$
$-\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{-is\epsilon_{\alpha}(k)}(h_{y})_{\alpha}(k)\bigg{)}e^{\tau B}\psi.$ (5.29)
Firstly, we show that the term $\mathcal{R}^{\lambda}_{\alpha}(k)\psi=\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi$ as defined in Lemma 5.1 constitutes indeed a small error. We estimate
$\sum_{l\in\Gamma}\sum_{\gamma\in\mathcal{I}_{l}}\|\mathcal{R}^{1}_{\gamma}(l)\psi\|^{2}$
$\leq\sum_{l\in\Gamma}\sum_{\gamma\in\mathcal{I}_{l}\cap\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma}|\eta_{\gamma}(k)|\int_{0}^{1}\textnormal{d}\tau\|\mathcal{E}_{\gamma}(k,l)e^{\tau B}\psi\|\bigg{)}^{2}$
@@
where we used $\sum_{\alpha\in\mathcal{I}_{k}}\bigg{(}\sum_{k\in\Gamma}|\eta_{\alpha}(k)|\bigg{)}^{2}\leq\sum_{k^{\prime},k\in\Gamma}\|\eta(k)\|_{l^{2}}\|\eta(k^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.6 and
@@
$\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\leq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\|c_{\alpha}^{\ast}(k)e^{(1-\tau)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|c_{\alpha}^{\ast}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
$\quad+\int_{0}^{1}\textnormal{d}\tau\sum_{k,l\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}|\eta_{\alpha}(l)|\ \|[c_{\alpha}^{\ast}(k),e^{(1-\tau)B}]\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|$
@@
$\quad+\sum_{k\in\Gamma}\|\eta_{s}\|\int_{0}^{1}\textnormal{d}\tau\bigg{(}\sum_{l\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|\mathcal{R}^{1}_{\alpha}(k)\mathcal{E}_{\alpha}(l,k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}$
@@
where we used $[c_{\alpha}^{\ast}(k),e^{\sigma B}]=-\lambda\eta_{\alpha}(k)e^{\sigma B}-e^{\sigma B}\mathcal{R}^{1}_{\alpha}(k)$ from Lemma 5.1, the Cauchy-Schwarz inequality for the $\alpha$-summation, Lemma A.3 in the third inequality and in the fourth inequality we used Proposition 5.4 and (5.30).
@@
$\|\text{Error}_{1}\|$ $\leq\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\left\{\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)+(1-e^{is\epsilon_{\alpha}(k)})\overline{(h_{y})_{\alpha}(k)}\right\}e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\leq Ck_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\quad+C\lambda\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}|\hat{V}(k)|\sum_{\alpha\in\mathcal{I}_{k}}\|n_{\alpha}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\leq Ck_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)e^{B}\mathcal{R}^{1}_{\alpha}(k)\psi\|$
$\quad+C\lambda k_{\textnormal{F}}\int_{0}^{t}\textnormal{d}s\|\hat{V}\|_{2}\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1}_{\alpha}(k)\psi\|^{2}\bigg{)}^{1/2}$
$\leq Ck_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}\int_{0}^{t}\textnormal{d}s\ (e^{C\|\eta_{s}\|}-1)\|(\mathcal{N}+3)^{2}\psi\|$
@@
$\leq C\ (f_{C}(\lambda k_{\textnormal{F}}t)-1)(\lambda+1)k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+3)^{2}\psi\|.$ (5.35)
where we used (A.2) and $e^{B}$ unitary in the third inequality and Lemma 5.6 in the last line. Using $\psi=\phi\otimes\Omega$ we obtain the desired bound.